OpenStudy (dr0zier99):

Which is the correct input-output table for the function f(x)= 3/x+4 ? Input (x) -3 -2 1 2 3 Output f(x) -1.5 -2 5 3.5 3 Input (x) -3 -2 1 2 3 Output f(x) 3 2.5 7 5.5 5 Input (x) -3 -2 1 2 3 Output f(x) -1 1.5 5 7 7.5 Input (x) -3 -2 1 2 3 Output f(x) 1 4 3.5 6 4.5

OpenStudy (anonymous):

Plug in the values for x, and see if they will give you the following values (output) when evaluated

OpenStudy (anonymous):

For example: \[f(-3) = \frac{ 3 }{ (-3)+4 } = \frac{ 3 }{ 1 } =1\]

OpenStudy (anonymous):

The -3 was your input

OpenStudy (anonymous):

Check all the inputs, if one is wrong then it's not that one.